OpenStudy (chelseatheruler):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1)

OpenStudy (campbell_st):

so to make things a little easier the given line in slope intercept form is y = -3x + 7 can you find the slope of the line..?

OpenStudy (chelseatheruler):

-3x?

OpenStudy (campbell_st):

not quite... its -3 so the slope of the perpendicular is the negative reciprocal.... so its 1/-(-3) = 1/3 so in slope intercept form the line is y = 1/3 x + b to find b, substitute your point values, x = 6 and y = -1 so you have -1 = 1/3 *(6) + b can you find the value of b..?

OpenStudy (chelseatheruler):

I really don't understand how to do this. :(

OpenStudy (campbell_st):

ok... so what is a third of 6

OpenStudy (chelseatheruler):

2

OpenStudy (chelseatheruler):

is that right?

OpenStudy (campbell_st):

thats correct so you have -1 = 2 + b whats the value of b

OpenStudy (chelseatheruler):

I dont know?

OpenStudy (campbell_st):

would 2 - 3 = -1...?

OpenStudy (anonymous):

The slope of the line. -3 you put in as B.

OpenStudy (chelseatheruler):

@campbell_st yes.

OpenStudy (campbell_st):

so the equation of your perpendicular line will be \[y = \frac{1}{3} x - 3\]

OpenStudy (chelseatheruler):

Thank you. :)